OpenStudy (anonymous):

Find the relative rate of change fro each of the following functions. f(x)=x^3 -3x^2

OpenStudy (anonymous):

I think you have to derive here, or am I wrong? Since it's about rate of change, right?

OpenStudy (anonymous):

no this is the problem

OpenStudy (anonymous):

the derivative would be 3x^2 -6x

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

that's how you compute the relative rate of change, by deriving ^_^

OpenStudy (anonymous):

im not understanding?

OpenStudy (anonymous):

so do i take f'(x) and divide it by f(x)

OpenStudy (anonymous):

lol, to find the relative rate of change, you have to find the derivative of the following function which is in this case :\[f(x) = x^3 - 3x^2\] so when you derive the function you'll get like what you've said : \[f'(x) = 3x^2 - 6x\] since there is no given for x, then that's the final equation I guess :)

OpenStudy (anonymous):

if they told you, for example, that x = 2 then substitute it in the derived function and you'll find the relative rate of change

OpenStudy (anonymous):

^_^ clearer?

OpenStudy (anonymous):

or am I mixing it up with something?

OpenStudy (anonymous):

i got all that, but my teacher still counted 2 points off?

OpenStudy (anonymous):

hmm, I'm missing a point aren't I?

OpenStudy (anonymous):

sstarica, when you're done here, can you help dinaortega? I'm leaving soon.

OpenStudy (anonymous):

I'll try my best loki :)

OpenStudy (anonymous):

could you help me lokisan?

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

relative rate of change is f'(x)/f(x)

OpenStudy (anonymous):

>_< thank you blexting

OpenStudy (anonymous):

blexting so how do i get that from f(x)= x^3 -3x^2 being the original problem and f'(x)=3x^2 -6x

OpenStudy (anonymous):

Sorry.. I don't know, but I knew the formula.. sstarica can you help?

OpenStudy (anonymous):

Post your assignment questions at www.aceyourcollegeclasses.com you can earn money answering questions there too!

OpenStudy (anonymous):

divide f'(x) with f(x)

OpenStudy (anonymous):

and simplify

OpenStudy (anonymous):

x^3-6x?

OpenStudy (anonymous):

6x/x^3

OpenStudy (anonymous):

\[\frac{f'(x)}{f(x)}=\frac{3x^2-6x}{x^3-3x^2}=\frac{3x-6}{x^2-3x}\]

OpenStudy (anonymous):

so you'll get :\[f'(x)/f(x) = (3x^2-6x)/(x^3-2x^2)\]\[= 3x(x-2)/x^2(x-2)\] \[= 3/x\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

sstarica.. the original was 3x^2. I agree with lokisan

OpenStudy (anonymous):

oh , sorry my bad ^^"

OpenStudy (anonymous):

is it clearer now ballards?

OpenStudy (anonymous):

thank you guys that makes since, but what if you have a log problem like this... f(x)=5x-xlnx

OpenStudy (anonymous):

it is clear how to figure out the relative rate of change. Could someone help me figure out the f'(x) of that log then i maybe could get it from there

OpenStudy (anonymous):

if f(x)=ln(x) f'(x)=1/x

OpenStudy (anonymous):

same story, find the derivative then divide it by the original function ^_^ and you'll get : f'(x) = 5 - [lnx + 1] ( use u'v + uv' to find the derivatice of xlnx) then divide it by the original : \[f'(x)/f(x) = (5-\ln|x|-1)/ 5x - xlnx\] and try to simplify :) give it a try now.